$SPAD/src/input r20bugs.input

Mike Dewar and Timothy Daly

December 3, 2016

Abstract

Contents

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool r20bugs.output
)set message test on
)set message auto off
)clear all

———-

– This file tests bugs fixed at release 2.0a. – Mike Dewar 17-5-95 — * —

)clear completely
--S 1 of 27
x := operator 'x
--R
--R
--R   (1)  x
--R                                                          Type: BasicOperator
--E 1
--S 2 of 27
sum( (x i - mu)**2, i=1..N )
--R
--R
--R         N
--R        --+       2                2
--R   (2)  >     x(i)  - 2mu x(i) + mu
--R        --+
--R        i= 1
--R                                                    Type: Expression(Integer)
--E 2
--S 3 of 27
D(%,mu)
--R

--R
--R         N
--R        --+
--R   (3)  >     - 2x(i) + 2mu
--R        --+
--R        i= 1
--R                                                    Type: Expression(Integer)
--E 3
)clear completely
--S 4 of 27
z := log(x+.3*%i)
--R
--R
--R   (1)  log(x + 0.3 %i)
--R                                             Type: Expression(Complex(Float))
--E 4
--S 5 of 27
z1 : EXPR Complex Float := z
--R
--R
--R   (2)  log(x + 0.3 %i)
--R                                             Type: Expression(Complex(Float))
--E 5
--S 6 of 27
complexForm(z1)$CTRIGMNP(Float, EXPR Complex Float)
--R
--R
--R                 2                0.3
--R   (3)  0.5 log(x  + 0.09) + atan(---)%i
--R                                   x
--R                                             Type: Complex(Expression(Float))
--E 6
)clear completely
--S 7 of 27
acot(-1)
--R
--R
--R        3%pi
--R   (1)  ----
--R          4
--R                                                    Type: Expression(Integer)
--E 7
)clear completely

--S 8 of 27
sqrt(-1.0::COMPLEX FLOAT)
--R
--R
--R   (1)  %i
--R                                                         Type: Complex(Float)
--E 8
--S 9 of 27
log(x+%i)::EXPR COMPLEX FLOAT
--R
--R
--R   (2)  log(x + %i)
--R                                             Type: Expression(Complex(Float))
--E 9
)clear completely
--S 10 of 27
positiveRemainder(-1,-5)
--R
--R
--R   (1)  4
--R                                                        Type: PositiveInteger
--E 10
)clear completely
--S 11 of 27
f:POLY FRAC COMPLEX INT := x^2+y^2
--R
--R
--R         2    2
--R   (1)  y  + x
--R                                 Type: Polynomial(Fraction(Complex(Integer)))
--E 11
--S 12 of 27
factor f
--R
--R
--R   (2)  (y - %i x)(y + %i x)
--R                       Type: Factored(Polynomial(Fraction(Complex(Integer))))
--E 12
)clear completely
--S 13 of 27
acot(-y)

--R
--R
--R   (1)  acot(- y)
--R                                                    Type: Expression(Integer)
--E 13
)clear completely
--S 14 of 27
a:=matrix [[1,2],[2,-1]]
--R
--R
--R        +1   2 +
--R   (1)  |      |
--R        +2  - 1+
--R                                                        Type: Matrix(Integer)
--E 14
--S 15 of 27
eigenMatrix a
--R
--R
--R        +   +-+       +-+    +
--R        |- \|5  + 1  \|5  + 1|
--R   (2)  |----------  --------|
--R        |     2          2   |
--R        |                    |
--R        +    1          1    +
--R                                 Type: Union(Matrix(Expression(Integer)),...)
--E 15
)clear completely
--S 16 of 27
positiveRemainder(-1::SINT,-5::SINT)
--R
--R
--R   (1)  4
--R                                                          Type: SingleInteger
--E 16
)clear completely

--   (1)  [[1.73046875 %i,- 2.0],[- 1.73046875 %i,- 2.0],[- 1.0,2.0],[1.0,2.0]]
--                                             Type: List(List(Complex(Float)))
--S 17 of 27 ok to fail. it seems there are 2 answers (see above)
complexRoots([u**2-v+1,v**2-4],[u,v],0.01)

--R
--R
--R   (1)
--R   [[1.732421875 %i,- 2.0],[- 1.732421875 %i,- 2.0],[- 1.0,2.0],[1.0,2.0]]
--R                                             Type: List(List(Complex(Float)))
--E 17
--S 18 of 27
complexRoots([u**2-v+1,v**2-4],[v,u],0.01)
--R
--R
--R   (2)  [[- 2.0,- 1.73046875 %i],[- 2.0,1.73046875 %i],[2.0,- 1.0],[2.0,1.0]]
--R                                             Type: List(List(Complex(Float)))
--E 18
)clear completely
--S 19 of 27
R := Record(key:Symbol,entry:Integer)
--R
--R
--R   (1)  Record(key: Symbol,entry: Integer)
--R                                                                 Type: Domain
--E 19
--S 20 of 27
tab := table([[a,1]$R,[b,2]$R,[c,3]$R])$Table(Symbol,Integer)
--R
--R
--R   (2)  table(c= 3,b= 2,a= 1)
--R                                                  Type: Table(Symbol,Integer)
--E 20
--S 21 of 27
remove!(b::Symbol,tab)
--R
--R
--R   (3)  2
--R                                                     Type: Union(Integer,...)
--E 21
)clear completely
--S 22 of 27
limit((x^a-a^x)/(x^x-a^a),x=a)
--R
--R
--R   (1)  0
--R                      Type: Union(OrderedCompletion(Expression(Integer)),...)
--E 22

)clear completely
--S 23 of 27
(4*x^6+16*x^4+16*x^2)*y^2+(-4*x^6+16*x^2)*y-8*x^6-16*x^4
--R
--R
--R           6      4      2  2        6      2       6      4
--R   (1)  (4x  + 16x  + 16x )y  + (- 4x  + 16x )y - 8x  - 16x
--R                                                    Type: Polynomial(Integer)
--E 23
--S 24 of 27
eval(%,x,z)
--R
--R
--R           2           6       2       4       2        2
--R   (2)  (4y  - 4y - 8)z  + (16y  - 16)z  + (16y  + 16y)z
--R                                                    Type: Polynomial(Integer)
--E 24
--S 25 of 27
exquo(%,expand factor %)
--R
--R
--R   (3)  1
--R                                         Type: Union(Polynomial(Integer),...)
--E 25
)clear completely
--S 26 of 27
log exp z
--R
--R
--R   (1)  z
--R                                                    Type: Expression(Integer)
--E 26
--S 27 of 27
normalize %
--R
--R
--R   (2)  z
--R                                                    Type: Expression(Integer)
--E 27
)spool
)lisp (bye)

———- 8

References^{[1] nothing}

9